OpenStudy (anonymous):

The sum of the circumference of 2 circles is 38pie cm. & the sum of their areas is 193pie cm^2. how do we calculate the radii of the circles?

jhonyy9 (jhonyy9):

the circumference formula is 2rpi and the area formula is pir2 where r2 signe r square

OpenStudy (anonymous):

sorry but what do we do first for the workings?

OpenStudy (anonymous):

C=2pi r x 2 (since there are 2 circles) 38pi = 4pi r 38pi/4pi=4pi/4pi r 38/4=r 9.5 =r A=pi r^2 x 2 (since there are 2 circles) 193pi=2pi r^2 193pi/2pi=2pi/2pi r^2 96.5= r^2 (96.5)^1/2=r 9.8=r that is only an approximation!! i hope i get it..

jhonyy9 (jhonyy9):

2pir1 +2pir2=38pi from this we result r1 +r2 =19 pi(r1)2 +pi(r2)2=193pi from thie result (r1)2 +(r2)2 =193

jhonyy9 (jhonyy9):

hope that after this will be easy to make the calcule

OpenStudy (anonymous):

i've checked the answer and the radii is 7cm and 12cm

jhonyy9 (jhonyy9):

- sign r1 = a r2 =b so than you get a+b=19 a2+b2=193 - where a2 is a square and b2 is b square

jhonyy9 (jhonyy9):

and now you need to solve this system for a and b

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

um so we get 2 equations.. a+b=19 a^2 (a squared)+b^2 (b squared)=193 ?

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

how do we continue from here?

jhonyy9 (jhonyy9):

from a+b=19 you get a=19-b so this substitute in the second equation and will get b=? after you will can calcule the a

OpenStudy (anonymous):

:o yeah. thanks jhonyy :D

jhonyy9 (jhonyy9):

my pleasure

jhonyy9 (jhonyy9):

thank you for the medal

OpenStudy (anonymous):

:) u deserve it haha.

jhonyy9 (jhonyy9):

ok good luck bye